OpenStudy (anonymous):

Write the equation y = -3x + 9 in standard form with integer coefficients.

OpenStudy (abb0t):

@dan815 @Kainui

OpenStudy (kainui):

@abb0t @dan815

OpenStudy (abb0t):

@dan815

OpenStudy (kainui):

@dan815

OpenStudy (superhelp101):

it is important to know the standard form which is ax + by + c = 0 y = -3x + 9 so then you put ^ this in standard

OpenStudy (texaschic101):

Ax + By = C is standard form....I thought

OpenStudy (superhelp101):

oh yeah oopps yep

OpenStudy (texaschic101):

so first add 3x to both sides of the equal sign and that should get you to your answer

OpenStudy (anonymous):

how do I add 3x 2 both sides @texaschic101

OpenStudy (texaschic101):

y = -3x + 9 ----add 3x to both sides 3x + y = -3x + 3x + 9 -- combine like terms 3x + y = 9 -- see how the -3x and the 3x cancel out on the right side....and this is now in standard form

OpenStudy (texaschic101):

or if it is easier to understand, write it like this... y = -3x + 9 + 3x + 3x ------------------add 3x + y = 9

OpenStudy (texaschic101):

just remember, what you do on one side of the equal sign, you have to do on the other side

OpenStudy (texaschic101):

any questions ?

OpenStudy (superhelp101):

nice;)

OpenStudy (texaschic101):

thanks....I try to explain

OpenStudy (anonymous):

@texaschic101 I have more ques on my test can u help meh

OpenStudy (texaschic101):

oh....a test.....um...not supposed to, might get me banned. Not supposed to, if its a test. I wish you wouldn't have said that.

OpenStudy (texaschic101):

Tell you what....close this question. Then ask me what you need help on. Even though the question is closed, you can still post stuff on it

OpenStudy (anonymous):

ok ill do that

OpenStudy (texaschic101):

ask away

OpenStudy (texaschic101):

??

OpenStudy (anonymous):

Referring to the figure, write the equation of the horizontal line through the point (0,3).

OpenStudy (anonymous):

ill use this same graph 4 tha next few ques

OpenStudy (texaschic101):

Learn this too... HOY -- horizontal line, 0 slope, represented by y = a number VUX -- vertical line, undefined slope, represented by x = a number A horizontal line will always have y = a number. In this case (0,3)...y = 3 and y will always equal 3. And your slope is 0. So your equation is : y = 0x + 3 (because the slope(m) = 0) or it can be written as y = 3

OpenStudy (anonymous):

Referring to the same graph, write the equation of the horizontal line through the point (3,0).

OpenStudy (texaschic101):

Horizontal line through (3,0) slope = 0 y = mx + b 0 = 0(3) + b 0 = b so your equation is : y = 0x + 0 or just y = 0...I think this one confuses me a little.

OpenStudy (anonymous):

Referring to the same graph, write the equation of the vertical line through the point (0,3).

OpenStudy (texaschic101):

VUX -- vertical line, undefined slope, x = a number The vertical line through (0,3)....x = 0 and y = 3 The x will always be 0. so your equation is : x = 0

OpenStudy (texaschic101):

if you have a vertical line, the the x's will always be the same number if you have a horizontal line, then the y's will always be the same number

OpenStudy (texaschic101):

and I know this is probably not gonna be used in your future, but you need to learn it to get out of high school

OpenStudy (anonymous):

Referring to the same graph, write the equation of the vertical line through the point (3,0).

OpenStudy (anonymous):

im sorry my internet was actin up :(

OpenStudy (texaschic101):

I am not sure about this one simply because it has an undefined slope, but I think it is just gonna be : x = 3 because (3,0)...x is 3

OpenStudy (anonymous):

thx 4 tha info @texaschic101 that info bout how 2 find out tha answer really helped thx :)

OpenStudy (texaschic101):

I try to help when I can. Good luck with the test

OpenStudy (anonymous):

i didn't do well :( tha peeps that helped meh b4 didn't help meh out enough i got a freakin D+ man @texaschic101

OpenStudy (anonymous):

gtg bye